OpenStudy (anonymous):

Tests by an independent auditing firm show that 50% of students in a certain school think that Algebra is the best course they have ever taken. Use the entire line of random numbers below to estimate the probability that a randomly selected group of five students will contain at least 3 Algebra lovers. Let the digits 1-5 represent a student who thinks that Algebra is the best course. 12024 55976 61475 70726 25408 62279 71874 03499 92659 26041 Convert your probability answer to a whole number representing a percent.

OpenStudy (anonymous):

so they are already in the groups of 5

OpenStudy (anonymous):

so there are 10 altogether

OpenStudy (anonymous):

I'm still confused..

OpenStudy (anonymous):

now find how many of them have 3 or more numbers that are less than 6 ( meaning they like algebra ) = 5 altogether

OpenStudy (anonymous):

the first group, third group , fifth, eigth and tenth

OpenStudy (anonymous):

probability = 5/10 = 50%

OpenStudy (anonymous):

The eighth group shouldn't count because zero doesn't count as a student who likes algebra (only digits 1-5). I get 4/10=40%